OpenStudy (noseboy908):

Write an equation of an ellipse in standard form with center at the origin and with the given vertex and co-vertex listed respectively. Vertex: (6,0) Co-Vertex: (0,-5)

OpenStudy (mertsj):

\[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2)}{b^2}=1\]

OpenStudy (mertsj):

That is the equation of an ellipse with center (h,k) and major axis horizontal. You also need to know that the distance from the center to the vertex is a and the distance from the center to the co-vertex is b

OpenStudy (noseboy908):

ok, so would it be: \[((x^2)/12^2)+((y^2)/10^2)\] ?

OpenStudy (noseboy908):

Or do I have the denominators switched?

OpenStudy (noseboy908):

yo still here @Mertsj ?

OpenStudy (perl):